5 people in a bus that has 7 stops left in it's journey. It's equally likely that any given passenger will leave at a given stop. What is the probability that no 2 passengers leave at the same stop.
Answer #1

probability that no 2 passengers leave at the same stop =P(all 5 people choose different stops out of 7)

=7P5 /75 =2520/16807